# An Adaptive Online Smoother with Closed-Form Solutions and Information-Theoretic Lag Selection for Conditional Gaussian Nonlinear Systems

### v1

@@ Appendix B Proof of the Optimal Online Forward-In-Time Discrete Smoother
Appendix C Details of the Online Expectation-Maximization Algorithm for Parameter Estimation
@@
In this appendix we outline the basic learning EM algorithm which sets the baseline structure for the online model identification algorithm developed in Section 4.3.1. As usual, it is assumed that the coupled CGNS in (2.1)–(2.2) is only partially observed, and based on the time discretization setup adopted thus far in this work (see (3.1)–(3.2) in Section 3.1), we denote the set of values of the state variables, when evaluated on the discrete-time steps $t_{j}$, as $\mathcal{X}=\{\mathbf{x}^{0},\ldots,\mathbf{x}^{j},\ldots,\mathbf{x}^{n}\}$ and $\mathcal{Y}=\{\mathbf{y}^{0},\ldots,\mathbf{y}^{j},\ldots,\mathbf{y}^{n}\}$, for the observable and unobservable components, respectively.
@@
$\ell(\boldsymbol{\theta}):=\log\left(p\big{(}\mathcal{X};\boldsymbol{\theta}\big{)}\right)=\log\left(\int_{\mathcal{Y}}p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathcal{Y}\right),$
where $\boldsymbol{\theta}$ is the collection of model parameters. This is also known as the marginal log-likelihood of the observed data, with the full log-likelihood being given by $\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right)$. Using now any distribution over the discretized hidden variables, with density $q(\mathcal{Y})$, a lower bound on the marginal log-likelihood can be obtained in a rather trivial manner in the following way [136]:
$\ell(\boldsymbol{\theta})=\log\left(\int_{\mathcal{Y}}p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}{\,\rm d}\mathcal{Y}\right)$ $=\log\left(\int_{\mathcal{Y}}q(\mathcal{Y})\frac{p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}}{q(\mathcal{Y})}{\,\rm d}\mathcal{Y}\right)$
[…]
$=\int_{\mathcal{Y}}q(\mathcal{Y})\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}-\int_{\mathcal{Y}}q(\mathcal{Y})\log(q(\mathcal{Y})){\,\rm d}\mathcal{Y}$
@@
where the negative value of $\int_{\mathcal{Y}}q(\mathcal{Y})\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}$ is the so-called free energy, while $\mathcal{S}(q)=-\int_{\mathcal{Y}}q(\mathcal{Y})\log(q(\mathcal{Y})){\,\rm d}\mathcal{Y}$ is the Shannon differential entropy. Therefore, based on the fact that $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ for all distributions $q$ over $\mathcal{Y}$, it is clear that maximizing the marginal log-likelihood is equivalent to maximizing $\mathcal{F}$ alternatively with respect to the distribution $q$ and the parameters $\boldsymbol{\theta}$ [142, 143], thus creating a bridge between the usual MLE algorithm and EM procedure. This alternate maximization of $\mathcal{F}$ is the essence of the EM algorithm, where the EM designation stems from the following procedure:
@@
[…]
@@
    $\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\left\{\mathcal{F}(q_{n+1},\boldsymbol{\theta})\right\}$
@@
An essential observation in this process, is that the maximization in the E-Step is exactly reached when $q$ is the posterior conditional distribution of $\mathcal{Y}$, that is
$q_{n+1}(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta}_{n}),$
@@
which in the online setting of sequential arrival of observations corresponds to the online smoother posterior distribution. This is attributed to the equality condition for the Jensen inequality, which we used to prove the bound between the marginal log-likelihood and the objective function $\mathcal{F}$. In Jensen’s integral inequality, $\phi(\mathbb{E}\left[X\right])\leq\mathbb{E}\left[\phi(X)\right]$ for $\phi$ being operator convex and $X$ an arbitrary-valued random variable, it is known that equality holds if and only if $\phi$ is an affine function or if $X$ is constant almost surely. As such, since the objective function $\mathcal{F}$ can be rewritten in the following much more intuitive form:
$\mathcal{F}(q,\boldsymbol{\theta})=-\mathcal{P}(q,p(\boldsymbol{\cdot}|\mathcal{X};\boldsymbol{\theta}))+\ell(\boldsymbol{\theta}),$
@@
where $\mathcal{P}$ is the relative entropy from Section 3.4.1, then equality holds when $\log\left(\frac{p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}}{q(\mathcal{Y})}\right)$ is an affine function, which is exactly achieved when $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})\propto p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}$, with the proportionality being due to Bayes’. Of course, this fact can also be seen through the Gibb’s inequality (see Section 3.4.1), since for $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})$ the relative entropy term on the right-hand side vanishes [144]. In such a situation, the bound $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ becomes an equality, $\mathcal{F}(q,\boldsymbol{\theta})=\ell(\boldsymbol{\theta})$. Note that the conditional distribution in the E-Step is very difficult to solve for general CTNDSs. Various numerical methods and approximations are often used [136, 135], which usually suffer from both approximation errors and the curse of dimensionality. Nevertheless, for the CGNS framework, the distribution $p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta}_{k})$ is given in an optimal and unbiased manner by the closed analytic formulae of the online discrete smoother from Theorem 3.1, which greatly facilitates the application of the online EM parameter estimation algorithm to many nonlinear models. On the other hand, since the differential entropy does not depend on $\boldsymbol{\theta}$, the maximum in the M-Step is obtained by maximizing the negative of the free energy,
@@
$\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\left\{\int_{\mathcal{Y}}p\big{(}\mathcal{Y}\big{|}\mathcal{X};\boldsymbol{\theta}_{n}\big{)}\log\left(p\big{(}\mathcal{X},\mathcal{Y};\boldsymbol{\theta}\big{)}\right){\,\rm d}\mathcal{Y}\right\}.$
@@
Finally, here we provide a simple numerical trick for accelerating the convergence of the parameters to their true values in the online EM algorithm. This process is based on momentum-based methods from classical convex optimization, which utilize momentum terms into the update to aid the learning process (e.g., Nesterov’s accelerated gradient descent [147]). We denote by $\boldsymbol{\theta}_{n}$ and $\boldsymbol{\theta}_{n+1}$ the learned parameters (or part of them) in the previous ($n$) and current ($n+1$) iteration, respectively. An acceleration of the parameter value at the current step can be achieved by
$\boldsymbol{\theta}_{n+1}^{\text{new}}=\boldsymbol{\theta}_{n}+\alpha\left(\boldsymbol{\theta}_{n+1}-\boldsymbol{\theta}_{n}\right),$
@@
where $\alpha\in(0,1]$ is a hyperparameter. If $\alpha=1$ then there is no acceleration; the acceleration rate depends on the amplitude of $\alpha$. Such a trick can be applied in the first few iterations of the online EM algorithm, especially for those in the unobserved process when the observability of the system is weak.

### v2

@@ Appendix B Proof of the Optimal Online Forward-In-Time Discrete Smoother
Appendix C Trade-off Analysis Between the Computational and Storage Complexity and Posterior Solution Skill for the Dyad Model in (4.1)–(4.2)
Figure C.1, similar to Figure 4.4, compares the computational time (in seconds) and storage requirements (in gigabytes) of the fixed- and adaptive-lag online smoothers against the NRMSE between their posterior mean time series and the true unobserved signal. As with Figure 4.4, computational times in Panels (b), (e), and (g) are averaged over multiple runs to reduce external fluctuations, with corresponding panels being plotted in a similar manner. Details on how the storage and time values are calculated are given in Appendix D. Comparing these with the computational behavior of the fixed- and adaptive-lag online smoothers in the LDA application (see Figure 4.4), which is a model that lacks the intermittent instabilities that emerge in the dyad model, (4.1)–(4.2), more intricate results arise.
Starting with the similarities between the two models, both the storage use (Panel (a), in red) and computational time (Panel (b), in red) of the fixed-lag online smoother show an algebraically increasing behavior with the fixed lag value. Furthermore, as is to be expected, since a constant upper lag bound parameter is used for all simulations of the adaptive-lag online smoother ($b\Delta t=3$ simulation time units), the storage use does not vary with decreasing $\delta$ (Panel (d), in red).
In contrast, the NRMSE of the fixed-lag method in the dyad model (Panels (a)–(b), in blue) roughly indicates an algebraic decrease with increasing fixed lag, instead of the exponential one in LDA (see Panels (a) and (b) of Figure 4.4). With regards to the adaptive-lag online smoother’s computational performance in time (Panel (e), in red), as well as its prediction skill formulated via the NRMSE (Panels (d)–(e), in blue), both are nontrivial functions of the tolerance parameter $\delta$. Specifically, with decreasing $\delta$, after a certain tolerance value, there is a regime switch from, roughly speaking, a sub-logarithmic increase for the former and algebraic to exponential decrease for the latter, to a logarithmic to algebraic increase for the former and algebraic decrease for the latter. In Panel (e), like the one in Figure 4.4, we also show at each data point on the computational time curve the percentage of time spent on the adaptive-lag calculation (including the calculation of the update values $D^{j,n-2}$). Due to the one-dimensional observable and unobservable state spaces, the time consumed on this operation is much smaller than that in the LDA case study, but, because of this, the storage need reductions are only marginal in nature. These observations collectively indicate how, in such settings, the adaptive-lag algorithm has the potential to provide a significant reduction in computational complexity beyond its fixed-lag counterpart, while maintaining a skillful state estimation in pathwise error statistics (even when using a relatively inflated upper lag bound of $b\Delta t=3$).
The aforementioned results also emerge in the recovery skill of these algorithms, formulated through the temporally-averaged information gain beyond the filter solution, depicted in Panels (c) for fixed lag and (f) for adaptive lag. In these, the associated signal and dispersion parts are also plotted. (Note how in the dyad model, due to the intermittent extreme events, the dispersion is comparatively much smaller than the signal part.) In the former, we observe how the time-averaged information gain algebraically increases with respect to the fixed lag value, while in the latter we see a somewhat sub-logarithmic behavior with decreasing $\delta$ (after the aforementioned tolerance parameter threshold is met).
Finally, in Panel (g), the results from using the LSDev sequence to define the adaptive lag (i.e., (3.30)–(3.31)) are plotted, similar to Panel (e). Since, as observed from Panel (h) in Figure 4.2, in this case study the extreme events do not allow for the standardized sequence of relative entropies or information gains in (4.3) to behave in an overall exponential manner in $j$ uniformly over $n$, as in the LDA numerical experiment (see Panel (b) and (d) in Figure 4.5), the LSDev approach showcases a slightly distinct behavior with respect to the tolerance parameter $\delta$ when compared to the approach which uses the original sequence (see Panel (e)). Specifically, there is a trade-off to be made; for larger NRMSE the use of the original sequence of relative entropies outperforms in terms of computational time, while for smaller ones, particularly for the apparent lower-bound limit at $0.6315$, the LSDev approach is able to terminate faster.
Figure C.1: Panels (a)–(f): Same as Panels (a)–(f) of Figure 4.4, but for the dyad-interaction model, (4.1)–(4.2), from the case study in Section 4.1. Panel (g): Same as Panel (e) but using the LSDev sequence to define the adaptive lags (per (3.30)–(3.31)).
Appendix D Details of the Storage Value and Computational Time Calculation in Figures 4.4 and C.1
The storage values (in gigabytes) depicted in Panels (a) and (d) of Figure 4.4 (and by extension in Panels (a) and (d) of Figure C.1) are calculated by adding the storage required for storing each of the MATLAB nested cell array structures that are used to hold the online smoother mean vectors $\boldsymbol{\mu}_{\text{s}}^{j,n}$, online smoother covariance matrices $\mathbf{R}_{\text{s}}^{j,n}$, and the online smoother update matrices $\mathbf{D}^{j,n-2}$ (one nested cell array for each), for that run of the online smoother algorithm (Algorithm 1). The outer cell array (of each nested cell array) is $1\times 1000$ in size, where $1000$ is the total number of observations in this numerical experiment. Each cell in the outer cell array holds another ($1\times n$)-sized cell array, where $n$ is the ordinal number of the current observation, which itself, in each of its cells, holds the associated multidimensional quantity mentioned prior (vector or matrix, respectively). Finally, all of these storage values also include the storage requirements for the 2D matrix holding the filter mean vectors and 3D matrices holding the filter covariance matrices and $\mathbf{E}^{j}$ and $\mathbf{F}^{j}$ auxiliary matrices (one for each). For the adaptive-lag online smoother we also include the 2D array holding the adopted adaptive-lag-defining criterion, $\left\{\mathcal{P}\left(p^{j,n}_{\text{updated}},p^{j,n}_{\text{lagged}}\right)\right\}_{R_{n}\leq j\leq n-1}$, or $\{\sigma^{j,n}\}_{R_{n}\leq j\leq n-1}$.
We utilize nested cell arrays for these case studies to simulate jagged or ragged 3D- and 4D-arrays, since the numerical experiments are not being carried out on an HPC with GPU clusters (and since MATLAB does not natively support jagged multidimensional arrays). Depending on the fixed lag value $L$ for the fixed-lag online smoother and the upper lag bound $b$ for the adaptive-lag online smoother, only the necessary components that are required to carry out the associated online smoother updates are calculated and stored, and therefore added to the aforementioned storage values. While this adopted approach leads to an overcalculation of the computational time required, as MATLAB can struggle with accessing elements in a (nested) cell array (due to how they occupy place in memory), this method allows us to run these algorithms efficiently even for high-dimensional systems without the need for an HPC with GPU clusters.
As for the computational time values, both for the fixed- and adaptive-lag algorithms, the simulation time does not include the time required for the calculation of the filter statistics, of $\mathbf{E}^{j}$ and $\mathbf{F}^{j}$ (see (3.5)–(3.6)), and of the model parameters at each newly obtained observation. Those are assumed to already be calculated and stored in memory for quick access. Therefore, the computational times depicted in these figures only cumulate the time needed at each $n$ (i.e., each observation) to access said memory and to carry out the the calculations in Algorithm 1 for the $j$’s defined either by the fixed or adaptive lag value, where for the latter we also include the time required to determine it (i.e., calculate the information gain sequence (and possibly its LSDev) and then use (3.32) (or (3.30))). Importantly, it does not include the time required to create and delete the aforementioned nested cell arrays that hold the online smoother Gaussian statistics and update matrices $\mathbf{D}^{j,n-2}$.
@@
Appendix E Details of the Online Expectation-Maximization Algorithm for Parameter Estimation
@@
In this appendix, we outline the basic learning EM algorithm, which sets the baseline structure for the online model identification algorithm developed in Section 4.3.1. As usual, it is assumed that the coupled CGNS in (2.1)–(2.2) is only partially observed, and based on the time discretization setup adopted thus far in this work (see (3.1)–(3.2) in Section 3.1), we denote the set of values of the state variables, when evaluated on the discrete-time steps $t_{j}$, as $\mathcal{X}=\{\mathbf{x}^{0},\ldots,\mathbf{x}^{j},\ldots,\mathbf{x}^{n}\}$ and $\mathcal{Y}=\{\mathbf{y}^{0},\ldots,\mathbf{y}^{j},\ldots,\mathbf{y}^{n}\}$, for the observable and unobservable components, respectively.
@@
$\ell(\boldsymbol{\theta}):=\log\big(p(\mathcal{X};\boldsymbol{\theta})\big)=\log\left(\int_{\mathcal{Y}}p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta}){\,\rm d}\mathcal{Y}\right),$
where $\boldsymbol{\theta}$ is the collection of model parameters. This is also known as the marginal log-likelihood of the observed data, with the full log-likelihood being given by $\log\big(p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})\big)$. Using any distribution over the discretized hidden variables, with density $q(\mathcal{Y})$, a lower bound on the marginal log-likelihood can be obtained in a rather trivial manner in the following way [105]:
$\ell(\boldsymbol{\theta})=\log\left(\int_{\mathcal{Y}}p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta}){\,\rm d}\mathcal{Y}\right)$ $=\log\left(\int_{\mathcal{Y}}q(\mathcal{Y})\frac{p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})}{q(\mathcal{Y})}{\,\rm d}\mathcal{Y}\right)$
$\geq\int_{\mathcal{Y}}q(\mathcal{Y})\log\left(\frac{p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})}{q(\mathcal{Y})}\right){\,\rm d}\mathcal{Y}\ (\text{by Jensen's integral inequality})$
$=\int_{\mathcal{Y}}q(\mathcal{Y})\log\big(p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})\big){\,\rm d}\mathcal{Y}-\int_{\mathcal{Y}}q(\mathcal{Y})\log\big(q(\mathcal{Y})\big){\,\rm d}\mathcal{Y}$
@@
where the negative value of $\int_{\mathcal{Y}}q(\mathcal{Y})\log\big(p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})\big){\,\rm d}\mathcal{Y}$ is the so-called free energy, while $\mathcal{S}(q)=-\int_{\mathcal{Y}}q(\mathcal{Y})\log\big(q(\mathcal{Y})\big){\,\rm d}\mathcal{Y}$ is the Shannon differential entropy. Therefore, based on the fact that $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ for all distributions $q$ over $\mathcal{Y}$, it is clear that maximizing the marginal log-likelihood is equivalent to maximizing $\mathcal{F}$ alternatively with respect to the distribution $q$ and the parameters $\boldsymbol{\theta}$ [111, 112], thus creating a bridge between the usual MLE algorithm and EM procedure. This alternate maximization of $\mathcal{F}$ is the essence of the EM algorithm, where the EM designation stems from the following procedure:
@@
    $q_{n+1}\leftarrow\underset{q}{\mathrm{argmax}}\big\{\mathcal{F}(q,\boldsymbol{\theta}_{n})\big\}$
@@
    $\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\big\{\mathcal{F}(q_{n+1},\boldsymbol{\theta})\big\}$
@@
An essential observation in this process is that the maximization in the E-Step is exactly reached when $q$ is the posterior conditional distribution of $\mathcal{Y}$, that is
$q_{n+1}(\mathcal{Y})=p\big(\mathcal{Y}\big|\mathcal{X};\boldsymbol{\theta}_{n}\big),$
@@
which in the online setting of sequential arrival of observations corresponds to the online smoother posterior distribution. This is attributed to the equality condition for the Jensen inequality, which we used to prove the bound between the marginal log-likelihood and the objective function $\mathcal{F}$. In Jensen’s integral inequality, $\phi\big(\mathbb{E}\left[X\right]\big)\leq\mathbb{E}\big[\phi(X)\big]$ for $\phi$ being a convex operator and $X$ an arbitrary-valued random variable, it is known that equality holds if and only if $\phi$ is an affine function or if $X$ is constant almost surely. As such, since the objective function $\mathcal{F}$ can be rewritten in the following much more intuitive form:
$\mathcal{F}(q,\boldsymbol{\theta})=-\mathcal{P}\big(q,p(\cdot|\mathcal{X};\boldsymbol{\theta})\big)+\ell(\boldsymbol{\theta}),$
@@
where $\mathcal{P}$ is the relative entropy from Section 3.4.1, then equality holds when $\log\left(\frac{p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})}{q(\mathcal{Y})}\right)$ is an affine function, which is exactly achieved when $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})\propto p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})$, with the proportionality being due to Bayes’. Of course, this fact can also be seen through the Gibbs’ inequality, since for $q(\mathcal{Y})=p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta})$ the relative entropy term on the right-hand side vanishes [113]. In such a situation, the bound $\mathcal{F}(q,\boldsymbol{\theta})\leq\ell(\boldsymbol{\theta})$ becomes an equality, $\mathcal{F}(q,\boldsymbol{\theta})=\ell(\boldsymbol{\theta})$. Note that the conditional distribution in the E-Step is very difficult to solve for general CTNDSs. Various numerical methods and approximations are often used [105, 104], which usually suffer from both approximation errors and the curse of dimensionality. Nevertheless, for the CGNS framework, the distribution $p(\mathcal{Y}|\mathcal{X};\boldsymbol{\theta}_{k})$ is given in an optimal and unbiased manner by the closed analytic formulae of the online discrete smoother from Theorem 3.1, which greatly facilitates the application of the online EM parameter estimation algorithm to many nonlinear models. On the other hand, since the differential entropy does not depend on $\boldsymbol{\theta}$, the maximum in the M-Step is obtained by maximizing the negative of the free energy,
@@
$\boldsymbol{\theta}_{n+1}\leftarrow\underset{\boldsymbol{\theta}}{\mathrm{argmax}}\left\{\int_{\mathcal{Y}}p\big(\mathcal{Y}\big|\mathcal{X};\boldsymbol{\theta}_{n}\big)\log\big(p(\mathcal{X},\mathcal{Y};\boldsymbol{\theta})\big){\,\rm d}\mathcal{Y}\right\}.$
@@
Finally, here we provide a simple numerical trick for accelerating the convergence of the parameters to their true values in the online EM algorithm. This process is based on momentum-based methods from classical convex optimization, which utilize momentum terms in the update to aid the learning process (e.g., Nesterov’s accelerated gradient descent [116]). We denote by $\boldsymbol{\theta}_{n}$ and $\boldsymbol{\theta}_{n+1}$ the learned parameters (or part of them) in the previous ($n$) and current ($n+1$) iteration, respectively. An acceleration of the parameter value at the current step can be achieved by
$\boldsymbol{\theta}_{n+1}^{\text{new}}=\boldsymbol{\theta}_{n+1}+\alpha\left(\boldsymbol{\theta}_{n+1}-\boldsymbol{\theta}_{n}\right),$
@@
where $\alpha\in[0,1]$ is a hyperparameter. If $\alpha=0$, then there is no acceleration; the acceleration rate depends on the amplitude of $\alpha$. Such a trick can be applied in the first few iterations of the online EM algorithm, especially for those in the unobserved process when the observability of the system is weak.
Appendix F Stability, Sensitivity, and Convergence of the Adaptive-Lag Online EM Parameter Estimation Algorithm for the Dyad Model (4.8)–(4.9) in Initial Parameter Values and Learning Duration
In this appendix we briefly discuss the stability, sensitivity, and convergence of the (adaptive-lag) online EM parameter estimation algorithm, when applied to the dyad model in (4.8)–(4.9), with respect to the algorithm’s burn-in or learning duration and initial parameter values.
The results noted and discussed in what follows have been derived on a purely empirical basis, through trial and error, due to the tremendous computational costs that a full parameter space analysis entails. Specifically, this empirical analysis is conducted in a “ceteris paribus” fashion, by adjusting only a single parameter’s initial guess or the algorithm’s learning period, while all other components of the algorithm remain unchanged from those in the simulation presented in Section 4.3.2. Furthermore, these adjustments are made in both directions, i.e., by both decreasing or increasing the component of interest, and done so to severe degrees as to confirm the algorithm’s stability or convergence skill is not localized under this regime.
For the dyad model in (4.8)–(4.9), the adaptive-lag online EM parameter estimation algorithm carried out to derive the results in Figure 4.6 is extremely stable, both in terms of the initial parameter values as well as in the length of the burn-in or learning period. Here, by stable, we mean that the algorithm does not destructively deteriorate and blow up; the algorithm might still diverge and approach a different subset of the parameter space as it evolves, but at least does so in a stable manner.
In terms of sensitivity and convergence skill with respect to the initial parameter values and burn-in or learning period, when we apply the adaptive-lag online EM parameter estimation algorithm on the dyad model in (4.8)–(4.9), we have:
  • $d_{u}$:
    Initial overestimation (i.e., a large positive value) affects $d_{u}$’s convergence, but it also affects that of $F_{v}$ (by overestimating it), as is to be expected; overdamping $u$ initially requires a larger positive $F_{v}$ to counteract this through the $\gamma u^{2}$ term. The convergence of the other parameters remains unaffected. Relatively large values of initial $d_{u}$ are needed to effectively diverge $d_{u}$ and $F_{v}$ away from their true values.
    If $d_{u}$ is set equal to zero initially, it remains close to its neighborhood and oscillates there even after extreme event observance. This is compensated by estimating a negative $F_{v}$ as the algorithm evolves. Other parameters’ convergence is again not severely affected by this.
    Initial underestimation (i.e., a large, in absolute value, negative initial guess) is not corrected with respect to the sign as the algorithm evolves, even after observing extreme events, with $d_{u}$ remaining at a significant negative value throughout. $F_{v}$ diverges as well, with large negative values. The convergence of the other parameters remains unaffected.
  • $\gamma$:
    Overestimating $\gamma$ initially severely skews the convergence of itself, $d_{v}$, and $F_{v}$, but for $\gamma$ and $F_{v}$ this is corrected with the observance of extreme events (especially after extreme event A; see Panel (a) in Figure 4.6). For $d_{v}$, it converges towards an extremely damped regime, where this is to be expected due to the $\gamma u^{2}$ term.
    By starting from zero, $\gamma$ remains close to it for the whole run, with $d_{u}$ being overestimated and the other parameters being underestimated.
    For negative initial $\gamma$, the algorithm’s estimation for it remains negative and converges towards $\gamma$, due to the symmetry of the energy conservation condition, but for large negative values it compensates by significantly overestimating $d_{v}$ and having a negative significant $F_{v}$. $d_{u}$ and $F_{u}$ are unaffected under this regime.
    Severe positive or negative $\gamma$ does slow down its convergence, but the true value is well approximated eventually after a sufficiently long run.
  • $F_{u}$:
    Significantly overestimating $F_{u}$ initially leads to the algorithm to search for an antidamping regime for $u$ over the parameter space (in $d_{u}$) and to severely underestimate $\gamma$, due to the quadratic coupling effects. It also completely misidentifies the $F_{v}$ value, due to the significant joint contribution of $F_{u}$ and $F_{v}$ in achieving system observability.
    Same for underestimating, with very large initial $F_{v}$ negative values, but in this case $\gamma$’s convergence is fine, with problems only emerging in the other parameters.
  • $d_{v}$:
    By initially severely overestimating $d_{v}$, the convergence can be severely skewed, leading to misidentification, because $v$ is the driving force behind the generation of the intermittent extreme events of the system (in $u$); after all, observing these instabilities accelerates and corrects the algorithm’s current estimation of the parameters.
    Same for underestimating, with very large negative values, which, although leading to an incorrect estimation for $d_{v}$, the algorithm at least immediately corrects the sign to a positive one (unlike in $d_{u}$’s case). The convergence of all other parameters is effectively unaffected.
    Starting at zero for $d_{v}$ initially has similar results as for negative initial values, i.e., they get quickly corrected, especially at the observance of extreme events.
    Finally, we note that even if $d_{v}$ is misidentified, as in the case of the results shown in Figure 4.6 of Section 4.3.2, leading to lower-order or time-series-based errors, nevertheless the memory and fidelity of the dyad model, represented by the higher-order metrics like the PDFs and ACFs of $u$ and $v$, are still effectively recovered.
  • $F_{v}$:
    Only for significant initial overestimation, it only slightly slows down and skews $d_{v}$’s convergence, while it marginally aids in the convergence of $d_{u}$ and $d_{v}$ when assumed to be negative and large.
  • $\begin{subarray}{c}\displaystyle\text{Learning}\\ \displaystyle\text{\qquad Period:}\end{subarray}$
    For the initial guess of $\boldsymbol{\theta}_{0}=(2,6,2,0.5,0.6)^{\mathtt{T}}$, we have the following effects from adjusting the length of the learning period from the $10$ time units used in Section 4.3.2 (in $[0,200]$).
    • Increase:
      Convergence skill, i.e., reaching the same values as the ones depicted in Figure 4.6 of Section 4.3.2, is not hindered by increasing the learning period. But, we do note that a really long burn-in period, combined with severely wrong initial parameter guesses (which are the ones utilized during it), can be rather destructive and lead to incorrect values and even severe instability and divergence. Also, a significantly long learning period can severely slow down convergence even for stable initial parameter values; e.g., by using a 30 time units-long learning period in this case, we miss out on doing state estimation at the observance of the highly influential extreme event A (see Panel (a) in Figure 4.6).
    • Decrease:
      Still stable, but trace plots are highly variable initially if chosen to be really brief. These variations are smoothed out relatively quickly, and satisfactory convergence is achieved at the observance of extreme events (but $d_{v}$ still converges to double the true value). Notably though, a training period that is at least a single time unit long is necessary for this case study; too small and $\gamma$ is thought to be zero, with similar results to as if we started with an initial guess of $\gamma=0$, as the learning period did not have enough data to showcase the presence of the energy-conserving quadratic coupling between $u$ and $v$.
We do note that it is possible to make this study into the sensitivity, stability and variability of the online EM parameter estimation algorithm much more rigorous, for this problem or model instance, by formulating an appropriate statistical response problem [78]: we identify the Fisher information matrix corresponding to the parameter estimation problem at hand as a function of a perturbation induced on the initial parameter values or learning period length, under sufficient regularity conditions, and then search for its most sensitive directions which are defined by its maximal eigenvectors (i.e., those corresponding to its largest eigenvalues).
Appendix G Results of the Fixed-Lag Online EM Parameter Estimation Algorithm on (4.8)–(4.9)
In Figure G.1, we showcase the trace plots of the parameter value absolute errors that are produced by the adaptive-lag and fixed-lag online EM parameter estimation algorithms (for the parameters of interest). By noting that the average adaptive lag, over $[10,200]$, of the adaptive-lag online EM algorithm is about $0.1686$ time units (see Panel (c) of Figure 4.6), for the fixed-lag online EM parameter estimation algorithm we use a constant lag of $L_{n}\Delta t=0.25$ time units for all $n$, i.e., for each newly acquired observation of the observed variable $u$. This amounts to about a $50\%$ increase from the mean adaptive lag.
As already observed in the adaptive-lag online EM algorithm implementation from Section 4.3.2, the observed extreme events help accelerate and correct the algorithm’s divergence, regardless of whether a fixed or adaptive lag is used. Additionally, as expected, these two variations of the online EM parameter estimation method do not exhibit significant discrepancies for the forcing parameters, $F_{u}$ and $F_{v}$, and the damping parameter of the unobserved variable, $d_{v}$. This is especially true after the algorithm is let run for long lead times.
However, it quickly becomes evident that the adaptive-lag variation is superior in correctly identifying the values of the parameters which are the driving force behind the significant dynamics of the model. Specifically, the adaptive-lag method is better at estimating $d_{u}$ and $\gamma$, and more importantly, the ratio $d_{u}/\gamma$; recall that whenever $v>d_{u}/\gamma$, this induces an instability in $u$, via anti-damping, which forms the system’s intermittent extreme events. This is especially apparent during the observation of extreme events, particularly after the highly influential events A and B. As such, despite the use of a fixed lag that is $50\%$ longer than the average one found in the adaptive-lag online EM algorithm, the fact that the latter adaptively chooses longer lags during the emergence of extreme events, thus capturing the significant contributions which influence the posterior smoother state estimation, aids in its efficacious model identification. Furthermore, by using negligible lags during periods of large signal-to-noise ratios (e.g., at the demise of the extreme events in $u$), instead of a uniformly large fixed lag value throughout the run, the adaptive-lag strategy can also comparatively show significant storage savings, depending of course on the choice of its upper lag bound parameter ($b$) and the fixed lag length (as is observed from the results in Figures 4.4 and C.1).
Figure G.1: Panels (a)–(f): Trace plots of the parameter value absolute errors produced by the adaptive-lag (in red) and fixed-lag (in blue) online EM parameter estimation algorithms.